OpenStudy (anonymous):

Solve the equation. 3(y + 6) = 30

OpenStudy (oaktree):

Divide by 3.

OpenStudy (oaktree):

y+6=10, and y=4.

OpenStudy (anonymous):

Divide By 3 !

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

3(y+6) = 30 3y+18 = 30 3y = 30-18 y = 12/3 =4